# The Statistical Fairness-Accuracy Frontier

### v1

@@ Abstract
Machine learning models must balance accuracy and fairness, but these goals often conflict, particularly when data come from multiple demographic groups. A useful tool for understanding this trade-off is the fairness–accuracy (FA) frontier, which characterizes the set of models that cannot be simultaneously improved in both fairness and accuracy. Prior analyses of the FA frontier provide a full characterization under the assumption of complete knowledge of population distributions—an unrealistic ideal. We study the FA frontier in the finite-sample regime, showing how it deviates from its population counterpart and quantifying the worst-case gap between them. In particular, we derive minimax-optimal estimators that depend on the designer’s knowledge of the covariate distribution. For each estimator, we characterize how finite-sample effects asymmetrically impact each group’s risk, and identify optimal sample allocation strategies. Our results transform the FA frontier from a theoretical construct into a practical tool for policymakers and practitioners who must often design algorithms with limited data.
@@ 1 Introduction
Across domains where predictive models guide consequential decisions—from lending and hiring to healthcare and education—society expects systems to be both accurate and fair. In supervised learning with multiple demographic groups, these objectives are often in tension: models that minimize overall error can yield unequal performance across groups, while interventions that reduce disparities may incur accuracy losses. A central, practically relevant challenge arises when a single predictive model must serve multiple groups, whether due to legal, logistical, or normative constraints. For instance, in the United States, the Equal Credit Opportunity Act (ECOA) prohibits lending decisions based on protected attributes such as race or sex and is commonly interpreted to disallow the use of different predictive models for different groups. Similarly, in hiring, automated screening systems typically apply a single predictive model to all applicants, since Title VII of the Civil Rights Act imposes legal constraints that, along with operational challenges, discourage the use of group-specific models (see Raghavan et al. [2020] for a discussion of U.S. employment discrimination law and the guidelines on “disparate treatment”).
In such cases—where group-specific models are disallowed or impractical—many existing models and algorithms are known to exhibit disparate performance across groups [Bartlett et al., 2022, Fuster et al., 2022, Dastin, 2022]; thus, understanding the best possible compromises between fairness and accuracy becomes critical. A principled way to approach this tension is through study of the fairness–accuracy (FA) frontier. The FA frontier consists of models for which it is impossible to improve fairness without sacrificing accuracy for at least one group, or to improve a group’s accuracy without harming fairness or the other group’s accuracy. This frontier allows decision-makers to visualize the explicit trade-offs and select a model that aligns with their fairness and performance objectives.
Recent work, notably by Liang et al. [2021], provides a valuable characterization of the FA frontier under the assumption that the population distributions for all groups are known exactly. In most real-world scenarios, however, this assumption is unlikely to hold—we are unlikely to possess such a distribution or to have access to the entire population. Instead, in practice, we must base our inferences and decisions regarding the FA frontier on a finite number of samples drawn from the underlying population. This paper tackles this finite-sample regime directly. We study how the empirical FA frontier deviates from its population counterpart, characterize the worst-case gap between them, and develop optimal estimators and sampling strategies. In short, we recast the FA frontier from an idealized object into a practical statistically-based mechanism design tool.
@@
Consider a two-group linear regression setting, with groups labeled “red” ($r$) and “blue” ($b$). For each individual in a group $g\in\{r,b\}$, the outcome variable is modeled as the inner product of the individual’s covariate vector and a group-specific parameter $\beta_{g}\in\mathbb{R}^{d}$, plus a mean-zero random noise term. The two groups may differ in both their underlying parameters, i.e., $\beta_{r}\neq\beta_{b}$, and in their covariate distributions. For any linear prediction model parameterized by $\beta$, we measure its group-specific accuracy as the risk—the expected squared loss under the distribution of group $g$. We measure fairness as the absolute difference in risk between the two groups, which is the equalized loss criterion [Zhang et al., 2019, Khalili et al., 2023]. Following Liang et al. [2021], we visualize the FA frontier in a two-dimensional plot, with the risks for the red and blue groups on the two axes. In this case, as shown in Figure 1(a), the FA frontier (under full knowledge of the distribution) traces a portion of the boundary of the set of all risk pairs (across the two groups) that are attainable by some linear model. Its two endpoints lie at the group-$r$ optimum $\beta_{r}$ and the group-$b$ optimum $\beta_{b}$.
@@
For homoskedastic linear regression, the FA frontier is group-balanced in the sense of Liang et al. [2021]: the most fair point lies between the two group-optimal endpoints. In this case, we can parametrize the FA frontier using a parameter $\lambda\in[0,1]$: for each $\lambda$, we define an objective function as a convex combination of group-specific risks, assigning weight $\lambda$ to the red group and $1-\lambda$ to the blue group. The pair of risks corresponding to the minimizer of this objective, denoted by $\beta_{\lambda}$, lies on the FA frontier and represents the first point of tangency between the frontier and a line of the form $\lambda x+(1-\lambda)y=c$ as $c>0$ increases, as illustrated in Figure 1(b). By varying $\lambda$ continuously from 0 to 1, we trace out the entire FA frontier.
@@
The choice of $\lambda$ is a central design decision shaped by the model’s legal and ethical context, reflecting the social planner’s preferences over group errors. To illustrate this, suppose 90% of the population belongs to the red group. A natural choice would then be to set $\lambda=0.9$. However, if we are training a predictive model for patients or for hiring across different demographics, we might want to ensure that the smaller group is not overshadowed. In that case, we could increase its weight and even assign equal importance to both groups, i.e., set $\lambda=0.5$. This approach, often referred to as “weighting by inverse class frequency,” is common in the fairness literature [Cui et al., 2019, Steininger et al., 2021]. There are also smoothed variants that interpolate between these extremes, offering a context-dependent balance between accuracy and fairness [Mikolov et al., 2013]. In summary, the choice of $\lambda$ ultimately rests with the designer, and the role of the FA frontier is to ensure that, whatever choice is made, the resulting algorithm is optimal for that setting.
@@
For subgaussian covariates satisfying the small-ball condition, we derive upper bounds for both the bias and variance terms. Our upper bound for the variance term naturally matches the bound in the known covariance case. For the bias term, we establish an upper bound that shows how it depends differently on the sample size, the covariance matrices, and the group heterogeneity $\|\beta_{r}-\beta_{b}\|$: it vanishes when the two parameters are equal and grows linearly with the squared norm of their difference. Moreover, for both terms, we establish matching lower bounds (up to constant factors), showing that the excess risk of any estimator must contain both the bias and variance terms, and that no estimator can improve upon the OLS estimator by more than a constant factor (see Theorem 2 and Theorem 3).
@@
Our bounds also provide an optimal sampling rule for a fixed total sampling budget, $n_{r}+n_{b}$. In the known-covariance case, we show that the optimal allocation of the sampling budget depends on the norm of the covariance matrix as well as the weight of that group’s risk in the overall loss ($\lambda$ versus $1-\lambda$), with the optimal sample size increasing as either quantity grows. This result is a direct analogue of Neyman’s classical allocation in stratified sampling [Neyman, 1934], where sample sizes are chosen in proportion to stratum variability and importance. The optimal sampling rule for the unknown covariance case is more nuanced. When the variance term dominates (e.g., when $\|\beta_{r}-\beta_{b}\|$ is small), the same allocation rule as in the known covariance case remains near-optimal. When heterogeneity is large and the bias term dominates, however, the optimal design shifts toward balancing the two sample sizes, $n_{r}\approx n_{b}$, to reduce covariance-estimation bias. Which regime governs depends on $\|\beta_{r}-\beta_{b}\|$, clarifying how problem structure should inform data collection when targeting a specific point on the FA frontier.
@@
Finally, while the preceding results bound the excess risk of the minimax-optimal empirical estimator with respect to the (weighted) average objective function, it is natural to ask: How does this estimator affect each group’s individual risk? As shown in Figure 2(b), we find that the empirical estimator does not affect the two groups equally. In particular, in the case of known covariance, we show that the increase in excess risk for the two groups is proportional to the ratio of the norms of their covariance matrix. In the unknown covariance case, beyond this difference arising from covariance scaling, the bias of the OLS estimator introduces an additional cross term. This term can shift the risks in opposite directions between groups and depends asymmetrically on $\lambda$ for the two groups. Interestingly, in both cases, the differing impact of the estimator on the two groups does not stem from differences in their sample sizes. These findings suggest that the fairness–accuracy trade-off implied by the empirical estimator may differ substantially in practice from the trade-off defined at the population level.
@@ 1.1 Related Work
We build on Liang et al. [2021], who formalize the FA Pareto frontier for assessing fairness-accuracy trade-offs under the assumption of full knowledge of the covariate and outcome distributions. Liu and Molinari [2024] relax the full-knowledge assumption and introduce a consistent estimator of the FA frontier, deriving the asymptotic distribution of the estimator and using it to design statistical tests for properties such as the gap between a given algorithm and the fairest alternative. Auerbach et al. [2024] further contribute a statistical test for determining whether a model achieving a given accuracy admits a Pareto improvement in fairness. We likewise address the practical challenges of analyzing the FA frontier from finite data, but take a fully non-asymptotic approach, providing finite-sample guarantees for estimators that target a social planner’s fairness-accuracy preferences.
@@
Our work contributes to the literature on fairness in regression, a topic that has traditionally received less attention than fair classification [see, e.g., Donini et al., 2018]. The most closely related work is that of Chzhen and Schreuder [2022], who develop a minimax-optimal fair regression estimator for an explicit fairness constraint based on the Wasserstein distance between the loss distributions across groups. Beyond this, several other works have introduced various notions of fairness for the regression setting and have proposed corresponding optimal estimation procedures [Calders et al., 2013, Berk et al., 2017, Fitzsimons et al., 2019, Agarwal et al., 2019, Chzhen et al., 2020a, b, Oneto et al., 2020] and for estimating the minimal utility cost associated with achieving fairness in regression [Zhao, 2021].
@@
A growing body of work bridges perspectives on fairness-accuracy trade-offs in computer science and statistics with analyses of social welfare [Hu and Chen, 2020, Liang and Lu, 2024, Gans, 2024, Rosenfeld and Xu, 2025]. Our results on the FA frontier can also be interpreted in this context if we view each individual’s utility as the risk associated with their group. More broadly, in the algorithmic fairness literature, it is widely recognized that no single fairness criterion can capture the diverse preferences that decision-makers may have over these trade-offs [Casacuberta et al., 2023, Verma and Rubin, 2018, Kleinberg et al., 2016]. Economists make this point in a similar way by documenting the heterogeneity in how individuals weigh equity against efficiency [Capozza and Srinivasan, 2024, Lockwood and Weinzierl, 2015]. Our framework builds on this connection by adopting a flexible, weight-based formulation that accommodates this kind of heterogeneity [Saez and Stantcheva, 2016, Hendren, 2020, Kasy and Abebe, 2021].
@@
Finally, much attention has been devoted to the fundamental question of when fairness and accuracy are in conflict and when they are aligned [Corbett-Davies et al., 2017, Wick et al., 2019, Blum and Stangl, 2019, Liang et al., 2021, Andersen and Maibom, 2020, Dutta et al., 2020, Cooper et al., 2021]. In settings where these objectives are at odds, a variety of methods have been developed to identify Pareto improvements, where either fairness or accuracy can be improved while keeping the other one unchanged. One line of work designs algorithms for computing fairer alternatives to a given model for specific applications [Fish et al., 2016, Little et al., 2022, Wei and Niethammer, 2022, Chohlas-Wood et al., 2024, Dehdashtian et al., 2024]. A closely related line of work draws on the legal notion of a Less Discriminatory Alternative, or a fairer model that maintains the original model’s accuracy [Gillis et al., 2024, Laufer et al., 2025, Samad et al., 2025]. Others take the reverse approach, focusing on accuracy improvements subject to fairness guarantees [Agarwal et al., 2018, Wang et al., 2023], or, more generally, methods for balancing multiple competing objectives in learning [Rolf et al., 2020, Martinez et al., 2020].
@@ 2 Framework
Following Liang et al. [2021], we measure the fairness of a prediction function $f$ using the risk disparity $|\mathcal{R}_{r}(f)-\mathcal{R}_{b}(f)|$, which quantifies the difference in performance of the model between the two groups. For regression tasks with a continuous outcome variable YY—the primary focus of this paper—this notion coincides with the definition of Equalized Loss proposed by Zhang et al. [2019]. [^1] For binary classifiers, i.e., when the outcome variable $Y$ is binary, Liang et al. [2021, Appendix O.1] provide a detailed discussion on how this definition relates to well-known fairness notions in the binary setting such as Equalized Odds [Hardt et al., 2016].

### v2

@@ Abstract
Machine learning models must balance accuracy and fairness, but these goals often conflict, particularly when data come from multiple demographic groups. A useful tool for understanding this trade-off is the fairness-accuracy (FA) frontier, which characterizes the set of models that cannot be improved in either fairness or accuracy without worsening the other. Prior analyses of the FA frontier provide a full characterization under the assumption of complete knowledge of population distributions—an unrealistic ideal. We study the FA frontier in the finite-sample regime, showing how it deviates from its population counterpart and quantifying the worst-case gap between them. In particular, we derive minimax-optimal estimators that depend on the designer’s knowledge of the covariate distribution. For each estimator, we characterize how finite-sample effects asymmetrically impact each group’s risk, and identify optimal sample allocation strategies. Our results transform the FA frontier from a theoretical construct into a practical tool for policymakers and practitioners who must often design algorithms with limited data.
@@ 1 Introduction
Across domains where predictive models guide consequential decisions—from lending and hiring to healthcare and education—society expects systems to be both accurate and fair. In supervised learning with multiple demographic groups, these objectives are often in tension: models that minimize overall error can yield unequal performance across groups, while interventions that reduce disparities may incur accuracy losses. A central, practically relevant challenge arises when a single predictive model must serve multiple groups, whether due to legal, logistical, or normative constraints. For instance, in the United States, the Equal Credit Opportunity Act (ECOA) prohibits lending decisions based on protected attributes such as race or sex and is commonly interpreted to disallow the use of different predictive models for different groups. Similarly, in hiring, automated screening systems typically apply a single predictive model to all applicants, since Title VII of the Civil Rights Act imposes legal constraints that, along with operational challenges, discourage the use of group-specific models (see raghavan2020mitigating for a discussion of U.S. employment discrimination law and the guidelines on “disparate treatment”).
In such cases—where group-specific models are disallowed or impractical—many existing models and algorithms are known to exhibit disparate performance across groups [bartlett2022consumer, fuster2022predictably, dastin2022amazon]. Thus, it becomes critical to characterize the best possible compromises between fairness and accuracy. A principled way to approach this tension is through study of the fairness–accuracy (FA) frontier. The FA frontier consists of models for which it is impossible to improve fairness without sacrificing accuracy for at least one group, or to improve a group’s accuracy without harming fairness or the other group’s accuracy. This frontier allows decision-makers to visualize the explicit trade-offs and select a model that aligns with their fairness and performance objectives.
Recent work, notably by liang2021algorithm, provides a valuable characterization of the FA frontier under the assumption that the population distributions for all groups are known exactly. In most real-world scenarios, however, this assumption is unlikely to hold—we are unlikely to possess such a distribution or to have access to the entire population. Instead, in practice, we must base our inferences and decisions regarding the FA frontier on a finite number of samples drawn from the underlying population. This paper tackles this finite-sample regime directly. We study how the empirical FA frontier deviates from its population counterpart, characterize the worst-case gap between them, and develop optimal estimators and sampling strategies. In short, we recast the FA frontier from an idealized object into a practical statistically-based mechanism design tool.
@@
Consider a two-group linear regression setting, with groups labeled “red” ($r$) and “blue” ($b$). For each individual in a group $g\in\{r,b\}$, the outcome variable is modeled as the inner product of the individual’s covariate vector and a group-specific parameter $\beta_{g}\in\mathbb{R}^{d}$, plus a mean-zero random noise term. The two groups may differ in both their underlying parameters, i.e., $\beta_{r}\neq\beta_{b}$, and in their covariate distributions. For any linear prediction model parameterized by $\beta$, we measure its group-specific accuracy as the risk—the expected squared loss under the distribution of group $g$. We measure fairness as the absolute difference in risk between the two groups, which is the equalized loss criterion [zhang2019group, khalili2023loss]. Following liang2021algorithm, we visualize the FA frontier in a two-dimensional plot, with the risks for the red and blue groups on the two axes. In this case, as shown in Figure 1(a), the FA frontier (under full knowledge of the distribution) traces a portion of the boundary of the set of all risk pairs (across the two groups) that are attainable by some linear model. For homoskedastic linear regression, the FA frontier is group-balanced in the sense of liang2021algorithm: the most fair point lies between the two group-optimal endpoints. Hence, as displayed in Figure 1(a), the two endpoints of the FA frontier lie at the group-$r$ optimum $\beta_{r}$ and the group-$b$ optimum $\beta_{b}$.
@@
In this case, we can parametrize the FA frontier using a parameter $\lambda\in[0,1]$: for each $\lambda$, we define an objective function as a convex combination of group-specific risks, assigning weight $\lambda$ to the red group and $1-\lambda$ to the blue group. The pair of risks corresponding to the minimizer of this objective, denoted by $\beta_{\lambda}$, lies on the FA frontier and represents the first point of tangency between the frontier and a line of the form $\lambda x+(1-\lambda)y=c$ as $c>0$ increases, as illustrated in Figure 1(b). By varying $\lambda$ continuously from 0 to 1, we trace out the entire FA frontier.
@@
The choice of $\lambda$ is a central design decision shaped by the model’s legal and ethical context, reflecting the social planner’s preferences over group errors. To illustrate this, suppose 90% of the population belongs to the red group. A natural choice would then be to set $\lambda=0.9$. However, if we are training a predictive model for patients or for hiring across different demographics, we might want to ensure that the smaller group is not overshadowed. In that case, we could increase its weight and even assign equal importance to both groups, i.e., set $\lambda=0.5$. This approach, often referred to as “weighting by inverse class frequency,” is common in the fairness literature [cui2019class, steininger2021density]. There are also smoothed variants that interpolate between these extremes, offering a context-dependent balance between accuracy and fairness [mikolov2013distributed]. In summary, the choice of $\lambda$ ultimately rests with the designer, and the role of the FA frontier is to ensure that, whatever choice is made, the resulting algorithm is optimal for that setting.
@@
For subgaussian covariates satisfying the small-ball condition, we derive upper bounds for both the bias and variance terms. Our upper bound for the variance term naturally matches the bound in the known covariance case. For the bias term, we establish an upper bound that shows how it depends differently on the sample size, the covariance matrices, and the group heterogeneity which is defined as $\|\beta_{r}-\beta_{b}\|$. Moreover, for both terms, we establish matching lower bounds (up to constant factors), showing that the excess risk of any estimator must contain both the bias and variance terms, and that no estimator can improve upon the OLS estimator by more than a constant factor (see Theorem 2 and Theorem 3).
@@
Our bounds also provide an optimal sampling rule for a fixed total sampling budget, $n_{r}+n_{b}$. In the known-covariance case, we show that the optimal allocation of the sampling budget depends on the norm of the covariance matrix as well as the weight of that group’s risk in the overall loss ($\lambda$ versus $1-\lambda$), with the optimal sample size increasing as either quantity grows. This result is a direct analogue of Neyman’s classical allocation in stratified sampling [neyman1934representative], where sample sizes are chosen in proportion to stratum variability and importance. The optimal sampling rule for the unknown covariance case is more nuanced. When the variance term dominates (e.g., when $\|\beta_{r}-\beta_{b}\|$ is small), the same allocation rule as in the known covariance case remains near-optimal. When heterogeneity is large and the bias term dominates, however, the optimal design shifts toward balancing the two sample sizes, $n_{r}\approx n_{b}$, to reduce covariance-estimation bias. Which regime governs depends on $\|\beta_{r}-\beta_{b}\|$, clarifying how problem structure should inform data collection when targeting a specific point on the FA frontier.
The results discussed thus far provide a complete characterization of how a given point on the FA frontier, parameterized by $\lambda$, is affected by the transition to the empirical setting. This characterization is particularly useful when the desired trade-off rate is known ex ante. However, policymakers may be interested in the entire FA frontier, as it illustrates the full range of feasible trade-offs. A key challenge in this case is that, because only a single dataset is available to derive all points on the frontier, the estimation errors are correlated across values of $\lambda$. We address this issue in Section 6 by deriving a uniform, high-probability bound on empirical error across the frontier; we simultaneously control how much the empirical pair of group risks deviates from the corresponding population risks on the frontier. This allows us to augment the empirical FA frontier with a high-probability confidence band for the entire frontier. We further show that the width of this confidence band decreases with sample size, being proportional to the maximum of the square roots of the inverse sample sizes across the two groups (see Theorem 4). This result provides a rigorous foundation for quantifying the uncertainty in policy decisions based on the empirical FA frontier.
@@
Finally, while the preceding results bound the excess risk of the minimax-optimal empirical estimator with respect to the (weighted) average objective function, it is natural to ask: How does this estimator affect each group’s individual risk? As shown in Figure 2(b), we find that the empirical estimator does not affect the two groups equally. In particular, when the covariance is known, we show that the ratio of the increases in excess risk across the two groups is proportional to the ratio of the norms of their covariance matrices. In the unknown covariance case, beyond this difference arising from covariance scaling, the bias of the OLS estimator introduces an additional cross term. This term can shift the risks in opposite directions between groups and depends asymmetrically on $\lambda$ for the two groups. Interestingly, in both cases, the differing impact of the estimator on the two groups does not stem from differences in their sample sizes. These findings suggest that the fairness–accuracy trade-off implied by the empirical estimator may differ substantially in practice from the trade-off defined at the population level.
@@ 1.1 Related Work
We build on liang2021algorithm, who formalize the FA Pareto frontier for assessing fairness-accuracy trade-offs under the assumption of full knowledge of the covariate and outcome distributions. liu2024inference relax the full-knowledge assumption and introduce a consistent estimator of the FA frontier, deriving the asymptotic distribution of the estimator and using it to design statistical tests for properties such as the gap between a given algorithm and the fairest alternative. auerbach2024testing further contribute a statistical test for determining whether a model achieving a given accuracy admits a Pareto improvement in fairness. We likewise address the practical challenges of analyzing the FA frontier from finite data, but take a fully non-asymptotic approach, providing finite-sample guarantees for estimators that target a social planner’s fairness-accuracy preferences.
@@
Our work contributes to the literature on fairness in regression, a topic that has traditionally received less attention than fair classification [see, e.g., donini2018empirical]. The most closely related work is that of chzhen2022minimax, who develop a minimax-optimal fair regression estimator for an explicit fairness constraint based on the Wasserstein distance between the loss distributions across groups. Beyond this, several other works have introduced various notions of fairness for the regression setting and have proposed corresponding optimal estimation procedures [calders2013controlling, berk2017convex, fitzsimons2019general, agarwal2019fair, chzhen2020fair, chzhen2020fairb, oneto2020general] and for estimating the minimal utility cost associated with achieving fairness in regression [zhao2021costs].
@@
A growing body of work bridges perspectives on fairness-accuracy trade-offs in computer science and statistics with analyses of social welfare [hu2020fair, liang2024algorithmic, gans2024algorithmic, rosenfeld2025machine]. Our results on the FA frontier can also be interpreted in this context if we view each individual’s utility as the risk associated with their group. More broadly, in the algorithmic fairness literature, it is widely recognized that no single fairness criterion can capture the diverse preferences that decision-makers may have over these trade-offs [casacuberta2023augmenting, verma2018fairness, kleinberg2016inherent]. Economists make this point in a similar way by documenting the heterogeneity in how individuals weigh equity against efficiency [capozza2024should, lockwood2015gustibus]. Our framework builds on this connection by adopting a flexible, weight-based formulation that accommodates this kind of heterogeneity [saez2016generalized, hendren2020measuring, kasy2021fairness].
@@
Finally, much attention has been devoted to the fundamental question of when fairness and accuracy are in conflict and when they are aligned [corbett2017algorithmic, wick2019unlocking, blum2019recovering, liang2021algorithm, andersen2020big, dutta2020there, cooper2021emergent]. In settings where these objectives are at odds, a variety of methods have been developed to identify Pareto improvements, where either fairness or accuracy can be improved while keeping the other one unchanged. One line of work designs algorithms for computing fairer alternatives to a given model for specific applications [fish2016confidence, coston2021characterizing, wei2022fairness, chohlas2024learning, dehdashtian2024utility]. A closely related line of work draws on the legal notion of a Less Discriminatory Alternative, or a fairer model that maintains the original model’s accuracy [gillis2024operationalizing, laufer2025constitutes, samad2025selecting]. Others take the reverse approach, focusing on accuracy improvements subject to fairness guarantees [agarwal2018reductions, wang2023aleatoric], or, more generally, methods for balancing multiple competing objectives in learning [rolf2020balancing, martinez2020minimax].
@@ 2 Framework
Following liang2021algorithm, we measure the fairness of a prediction function $f$ using the risk disparity $|\mathcal{R}_{r}(f)-\mathcal{R}_{b}(f)|$, which quantifies the difference in performance of the model between the two groups. For regression tasks with a continuous outcome variable YY—the primary focus of this paper—this notion coincides with the definition of Equalized Loss proposed by zhang2019group. [^1] For binary classifiers, i.e., when the outcome variable $Y$ is binary, liang2021algorithm provide a detailed discussion on how this definition relates to well-known fairness notions in the binary setting such as Equalized Odds [hardt2016equality].